OpenStudy (anonymous):

PLEASE HELP If the following is a polynomial function, then state its degree and leading coefficient. If it is not, then state this fact. f(x) = -16x5 - 7x4 - 6

OpenStudy (campbell_st):

in your question the leading term is \[-16x^5\] it is the term that contains the highest power... the degree is the power of the leading term. the coefficient is the number associated with x hope that helps

OpenStudy (anonymous):

can it even be a polynomial function if the sixteen is negative?

OpenStudy (campbell_st):

it can think about the quadratic \[y=x^2~~~ and ~~~y=-x^2\] the negative causes a reflection in the x axis... same applies to this

OpenStudy (anonymous):

okay so the degree is 5 and the leading term is -16?

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

cool! can you help me with one more problem?

OpenStudy (campbell_st):

ok...

OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (anonymous):

i know that you do like x-2 and x+4 but i dont know what to do with the imaginary numbers

OpenStudy (campbell_st):

ok.... so the easy bit x = 2 is a zero then (x -2) is a factor so you can find the binomial factor for x = -4

OpenStudy (campbell_st):

ok... so then the complex roots there 2 of those \[x = 1 + 3i~~~~and ~~~~x = 1 - 3i\] so handle them is a similar way subtract 1 from both sides x - 1 = 3i x - 1 = -3i square both sides of the equation \[(x - 1)^2 = 9i^2~~~~~and ~~~~~ (x -1)^2 = 9i^2 \] remember (-3)^2 = 9 so now you need to know that i^2 = -1 so substituting you get \[(x -1)^2 = -9~~~~or~~~~~(x - 1)^2 + 9 = 0\] I'd expand and simplify it to a quadratic factor of \[(x^2 - 2x + 10)\] now you just need to multiply the 3 factors for the equation

OpenStudy (anonymous):

so foil them all??

OpenStudy (anonymous):

i got x^4-2x^2+4x-80

OpenStudy (campbell_st):

yes... so there really isn't an easy way other than just writing them in factored form... I suppose it depends on how the question is written.

OpenStudy (anonymous):

but its not an answer choice? did i do something wrong?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

i found my error it should be x^4-2x^2+36x-80

OpenStudy (campbell_st):

well I think you are missing a term in x^3 (x^2 - 2x + 10)(x^2 + 2x - 8) so Id do x^2(x^2 -2x + 10) + 2x(x^2 -2x + 10) - 8(x^2 -2x + 10) and then collect like terms

OpenStudy (anonymous):

okay thank you very much!